# Instability of a fluctuating biomimetic membrane driven by an applied uniform DC electric field

### v1

@@ Abstract
The linear stability of a lipid membrane under a DC electric field, applied perpendicularly to the interface, is investigated in the electrokinetic framework, taking account to the dynamics of the Debye layers formed near the membrane. The perturbed charge in the Debye layer redistributes and destabilizes the membrane via electrical surface stress interior and exterior to the membrane. The instability is suppressed as the difference in the electrolyte concentration of the solutions separated by the membrane increases, due to a weakened base state electric field near the membrane. This result contrasts with the destabilizing effect predicted using the leaky dielectric model in cases of asymmetric conductivity. We attribute this difference to the varying assumptions about the perturbation amplitude relative to the Debye length, which result in different regimes of validity for the linear stability analysis within these two frameworks.
@@ I Introduction
Cells are enveloped by thin ion-impermeable membranes that isolate the cell interior form the environment. The response of biological membranes and cells to electric fields has received a lot of attention, due to practical applications in cell manipulation, such as fusion and electroporation [1, 2]. These processes rely on the controlled destabilization of membranes [3, 4], where electric fields influence membrane fluctuations, alter geometry through stretching and compression [5, 6], and modify mechanical properties such as tension and bending [7]. Despite extensive experimental [5, 6] and theoretical studies, the fundamental mechanisms governing membrane destabilization in electric fields remain an area of ongoing research. Theoretical models for instability analysis involve two main approaches: the leaky dielectric model (LDM) and the Poisson-Nernst-Planck (PNP) model.
@@
The leaky dielectric model (LDM) assumes that free charge is confined to interfaces separating media with different electric properties (conductivity and permittivity), while the bulk region is charge-free. Pioneering work in this area dates back to the 1960s, with notable contributions from G. I. Taylor and J. R. Melcher [8, 9], who investigated the deformation and stability of fluid-fluid interfaces under electric fields. The LDM was extended to model electrohydrodynamics of bio-membranes by treating the thin membrane as a zero-thickness interface with effective electric properties, e.g., capacitance $C_{m}$. Studies using this model have reported instabilities induced by various mechanisms, including ion currents across the membrane subjected to DC normal fields [10] and tangent fields [11], and capacitive current upond application of a DC field [12], and in AC fields with periods shorter than the capacitor charging time [13].
@@
Although effective on a theoretical basis, the above studies neglect the thermal motion of the free charge. In a real system, the ions near an interface form a diffuse layer. This layer has a characteristic thickness, called the Debye screening length, which is typically few nanometers. The Poisson-Nernst-Planck equation governs the dynamics of the charges within electrolytic materials, and it has been used extensively in electrokinetics to study membrane or fluid interfaces [14]. Ref. [15] investigated curvature-surface charge density coupling induced instabilities of the membrane, where the charges can flip to the other side and where they can diffuse on the same side, and [16] focused on such instability with strongly charged surfaces. [17, 18] analyzed the electrostatic correction to the surface tension bending rigidity on the instability with a finite-thickness membrane. [19, 20] proposed a zero-thickness model with Robin-type of boundary conditions, predicting electrokinetic corrections to the elastic moduli of the membrane, and a new destabilizing term in the growth rate of the membrane undulations due to ion current through the membrane. These studies report membrane instability arising from unbalanced electrical stress on the membrane, due to either surface charge [15, 16] or charge within the Debye layer [17, 18, 19, 20]. Similar instabilities have also been observed in fluid-fluid interface systems, such as the velocity enhancement at electrolyte/liquid metal interfaces [21], and the perfect-dielectric liquid/electrolyte interfaces [22].
@@
In a steady DC electric field normal to a non-conducting membrane, the LDM predicts that the fully-charged membrane capacitor suppresses instabilities [10]. Instability arises only in AC fields [13] or during the startup of a DC field [12], where the membrane acts as a short-circuited capacitor, leading to charge imbalances on its surfaces. However, the PNP model identifies a different mechanism, predicting instability even under steady-state DC fields due to negative tension in the membrane [19, 20]. The physical origin of this negative tension remains less intuitive, and we provide an explanation of the instability mechanism by examining charge redistribution within the Debye layer. Ref.[19, 20] focused on symmetric membranes, i.e. separating electrolyte solution with the same ions and concentrations. We extend the analysis to asymmetric electrolyte concentrations and demonstrate that such asymmetry has a stabilizing effect, mitigating the instability predicted by the PNP model. This highlights the critical role of ionic distribution and membrane asymmetry in determining stability under electric fields. In Section II, we introduce the governing equations for two electrolyte layers of ions, separated by a thin ion-impermeable membrane. Section III discusses the base state solution and its linear stability. We further explore the effects of key parameters on instability in Section IV, identifying the sources of instability and elucidating why conductivity asymmetry reduces it. In addition, we show such asymmetric conductivity will result into the spontaneous curvature. Finally, conclusions and avenues for further research are summarized in Section V.
@@ II Problem formulation and Governing equations
The system consists of two electrolyte solutions separated by a thin planar membrane impermeable to ions as shown in (Fig. 1). The concentrations of the positive and negative ions are $n^{+}_{i}$ and $n^{-}_{i}$, where $i=1,2$ refer to the upper fluid and the lower fluid, respectively. The electrodes are located at $z=\pm L$, where DC voltage of $2V$ is applied. The position of the flat membrane is $z=0$. We consider the two electrolytes are both $1:1$, with the same permittivity $\varepsilon_{f}$, but different concentration $n_{i}^{\infty}$. This implies a different conductivity of the electrolyte solutions, $\sigma_{i}=2e^{2}\omega n_{i}^{\infty}$ [14], where $\omega$ is the mobility of ions and $e$ is the elementary charge. We consider the membrane to be impermeable to ions.
Figure 1: A planar lipid bilayer membrane, with shape denoted by $z=h(x)$, separating electrolyte solutions with different concentration. The electrodes, with potential $\phi=\pm V$, are held at $z=\pm L$.
@@ II.1 Governing equations
where $\rho_{i}=e(n_{i}^{+}-n_{i}^{-})$ is the charge density, which obeys the linearized Poisson-Nernst-Planck equations
@@
The perturbations in the concentration $n_{i}^{\pm}=n_{i}^{\infty}+\delta n_{i}^{\pm}$ [20] are used to linearize the PNP, assuming small applied voltages compared to thermal voltage $k_{B}T/e$ [23]. The $\bm{j}_{i}$ is the electric current density, $k_{B}T$ is the thermal energy, $D$ is the ion diffusion coefficient (all ions are assumed to have the same mobility), and $\bm{v}_{i}$ is the fluid velocity. Here we introduce the Debye length
$\lambda_{D_{i}}=\left(\frac{2e^{2}n^{\infty}_{i}}{\varepsilon_{f}k_{B}T}\right)^{-1/2}$ (3)
that defines the characteristic length scale for charge relaxation in the electrolyte, and we define $\lambda_{D}=(\lambda_{D_{1}}\lambda_{D_{2}})^{1/2}$ as the composite Debye length in our asymmetric system, following the notation of [24].
@@
where $\mu_{i}$ is the fluid viscosity and $-\rho_{i}\nabla\phi_{i}$ is the Coulomb body forces due to the electric field acting on the charge distributed in the bluk.
@@ II.2 Boundary conditions
At the top and bottom electrodes $z=\pm L$, we impose the voltage $\phi_{i}(z=\pm L)=\pm V$, which neglects the polarization at the surface at electrodes [20] since we focus on the membrane dynamics. The distance between the electrodes is assumed to be much larger than the Debye length $L\gg\lambda_{D}$, such that the bulk electrolyte is neutral and charge density is negligible $\rho_{i}(z=\pm L)=0$.
The membrane’s position is represented by $z=h(x)$, with unit normal vector $\bm{n}=(-h_{x},1)/(1+h_{x}^{2})^{1/2}$ pointing upward. On the membrane, we apply a Robin-type boundary condition [19, 20]
@@
[…]
since $z=+\infty$ is taken as the position far from the membrane and electric field vanishes in the bulk $-\frac{\partial\phi_{1}}{\partial z}|_{z=+\infty}=0$. The macroscopic charge on the other side of the membrane, denoted as $Q_{2}$, can be calculated similarly. Upon analysis, the Robin BC (5) further suggests $Q_{1}=-Q_{2}$, signifying that the total charge on both sides of the membrane always remains balanced, regardless of conductivity.
Another interfacial condition is the non-flux condition [25]:
@@
This indicates an insulating membrane, i.e., charge is not permitted to cross the membrane.
@@
where we denote ${[\![}f{]\!]}_{a}=f_{1}(z=a)-f_{2}(z=a)$ as the jump of $f$ at position $z=a$, and $\bm{T}$ as the total stress consists with the hydrodynamic stress ${\bm{T}^{hd}}$ and Maxwell’s stress ${\bm{T}^{el}}$
@@
where $H$ is the curvature, $K_{G}$ is the Gaussian curvature, and $\Gamma$ is the surface tension. $C_{0}$ is the bilayer spontaneous curvature arising from the asymmetry of the monolayers. In this stability study, we will initially neglect $C_{0}$ by assuming $C_{0}d_{0}\ll 1$ [27] for a membrane of zero thickness. We will address its effects in Sec. IV.3.
@@
$v=\frac{\partial h}{\partial t}+u\frac{\partial h}{\partial x}.$ (11)
@@ II.3 Dimensionless parameters and rescaling
The parameters in the above equations are scaled using their characteristic values in the system:
@@
$(x,z)$ $\mapsto$ $\lambda_{D}(x,z),\quad h\mapsto\lambda_{D}h,\quad d_{0}=\lambda_{D}d,$ (12)
$t$ $\mapsto$ $\underbrace{\frac{\mu\lambda_{D}^{2}}{\varepsilon_{f}\phi_{c}^{2}}}_{t_{el}}t,\quad p\mapsto\frac{\mu}{t_{c}}p,$
where $\phi_{c}$ is the thermal voltage, and $t_{el}$ is the time scale in electrohydrodynamic flow [10]. Using this time scale, the scaled PNP model is expressed as:
@@
where $R=\frac{\sigma_{1}}{\sigma_{2}}=\frac{n_{1}^{\infty}}{n_{2}^{\infty}}$ denotes the conductivity ratio. With $R\mapsto R^{-1}$, Eqs. (13) applies for phase 2. The $\alpha$ represents the time scale ratio between $t_{el}$ and the characteristic charging time of the double-layer capacitor $t_{c}=\frac{\lambda_{D}L}{D}$ [23, 24]:
@@
In this study, we consider the situation where $\alpha\ll 1$, and set $\alpha=0$. In this limit, the charge relaxation in Debye layer happens instantaneously, and the Debye layer remains at equilibrium when subjected to perturbations. The dimensionless Robin BC is given by
@@
is the dimensionless capacitance. The non-flux condition at the membrane remains the form in Eq. (7) with the dimensionless current
@@
${[\![}-p+2\partial_{z}v+\frac{1}{2}(\partial_{z}\phi)^{2}-\frac{1}{2}\beta d(\partial_{z}\phi_{m})^{2}{]\!]}_{z=h}\\ =-\frac{1}{Ca}[4H^{3}-4HK_{G}+2\nabla_{s}^{2}H-2\bar{\Gamma}H],$ (19)
where the capillary number $Ca$ compares the electric and bending stresses:
@@
The electric traction at the membrane consists of two parts, with ${[\![}\frac{1}{2}(\partial_{z}\phi)^{2}{]\!]}$ coming from the electric field in the external medium, and ${[\![}-\frac{1}{2}\beta d(\partial_{z}\phi_{m})^{2}{]\!]}$ coming from the electric field $\bm{E}_{m}=-\nabla\phi_{m}$ in the membrane [28, 29]. We emphasize that although the Robin condition (15) effectively maps the electric field and charge density on both side of the membrane to a zero-thickness membrane, the $\phi_{m}$ term in Eq. (19) remains necessary to calculate the effective stress in the zero-thickness limit [17, 20, 18]. We will address it in Sec. III.

### v2

@@ Abstract
The linear stability of a lipid membrane under a DC electric field, applied perpendicularly to the interface, is investigated in the electrokinetic framework, taking into account the dynamics of the Debye layers formed near the membrane. The perturbed charge in the Debye layers redistributes and generates destabilizing Maxwell stress on the membrane, which outweighs the stabilizing contribution from the electrical body force, leading to a net destabilizing effect. The instability is suppressed as the difference in the electrolyte concentration of the solutions separated by the membrane increases, due to a weakened base state electric field near the membrane. This result contrasts with the destabilizing effect predicted using the leaky dielectric model in cases of asymmetric conductivity. We attribute this difference to the varying assumptions about the perturbation amplitude relative to the Debye length, which result in different regimes of validity for the linear stability analysis within these two frameworks.
@@ I Introduction
Cells are enveloped by thin membranes that isolate the cell interior from the environment. The response of biological membranes and cells to electric fields has received a lot of attention, due to practical applications in cell manipulation, such as fusion and electroporation [1, 2]. These processes rely on the controlled destabilization of membranes [3, 4], where electric fields influence membrane fluctuations, alter geometry through stretching and compression [5, 6], and modify mechanical properties such as tension and bending [7]. Despite extensive experimental [5, 6] and theoretical studies, the fundamental mechanisms governing membrane destabilization in electric fields remain an area of ongoing research. Theoretical models for instability analysis involve two main approaches: the leaky dielectric model (LDM) and the Poisson-Nernst-Planck (PNP) model.
@@
The LDM assumes that free charge is confined to interfaces separating media with different electric properties (conductivity and permittivity), while the bulk region is charge-free. Pioneering work in this area dates back to the 1960s, with notable contributions from G. I. Taylor and J. R. Melcher [8, 9], who investigated the deformation and stability of fluid-fluid interfaces under electric fields. The LDM was extended to model electrohydrodynamics of bio-membranes by treating the thin membrane as a zero-thickness interface with effective electric properties, e.g., capacitance $C_{m}$. Studies using this model have reported instabilities induced by various mechanisms, including ion currents across the membrane subjected to DC normal fields [10] and tangent fields [11], and capacitive current upon application of a DC field [12], and in AC fields with periods shorter than the capacitor charging time [13].
@@
Although effective on a theoretical basis, the above studies neglect the thermal motion of the free charges. In a real system, the ions near an interface form a diffuse layer. This layer has a characteristic thickness, called the Debye screening length, which is typically a few nanometers. The PNP equations govern the dynamics of the charges within electrolytic materials, and have been used extensively in electrokinetics (EK) to study membrane or fluid interfaces [14]. Refs. [15, 16] investigated curvature-surface charge density coupling induced instabilities of a membrane. Refs. [17, 18] analyzed the electrostatic correction to the tension and bending rigidity, and the instability of a finite-thickness membrane. Refs. [19, 20] proposed a zero-thickness model with the Robin-type boundary condition, predicting electrokinetic corrections to the elastic moduli of the membrane, and a new destabilizing term in the growth rate of the membrane undulations due to ion current through the membrane. These studies report membrane instability arising from unbalanced electrical stress on the membrane, due to either surface charge [15, 16] or charge within the Debye layer [17, 18, 19, 20]. Similar instabilities have also been observed in fluid-fluid interface systems, such as the perfect-dielectric liquid/electrolyte interfaces [21].
@@
In a steady DC electric field normal to a non-conducting membrane, the LDM predicts that the fully-charged membrane capacitor suppresses instabilities [10]. Instability arises only in AC fields [13] or during the startup of a DC field [12], where the membrane acts as a short-circuited capacitor, leading to charge imbalances on its surfaces. However, the PNP model predictis instability even under DC fields [19, 20]. The physical mechanisms of this instability remain less intuitive and we provide an explanation by examining the charge distribution in the Debye layers near the membrane. Refs. [19, 20] focused on symmetric membranes, i.e., separating electrolyte solutions with the same ions and concentrations. However, in the experiments with vesicles in electric fields[22, 23, 24, 25, 26], the electrolyte concentrations inside and outside the vesicle are typically different, necessitating an analysis of asymmetric membranes. We extend the analysis to account for this asymmetry and show that differences in ionic concentration can have a stabilizing effect, reducing the instability predicted by the symmetric model. Our analysis highlights the critical role of ionic distribution and membrane asymmetry in determining stability under electric fields.
In Section II, we introduce the governing equations for two electrolyte solutions of ions, separated by a thin ion-impermeable membrane. Section III discusses the base state solution and its linear stability. We explore the effects of key parameters on the instability in Section IV, identifying the sources of instability and elucidating why conductivity asymmetry reduces it. In addition, we show that asymmetric conductivity gives rise to the spontaneous curvature. Finally, conclusions and avenues for further research are summarized in Section V.
@@ II Problem formulation and Governing equations
The system consists of two electrolyte solutions separated by a thin planar membrane impermeable to ions, as shown in Fig. 1. The concentrations of the positive and negative ions are $n^{+}_{i}$ and $n^{-}_{i}$, where $i=1,2$ refers to the upper fluid and the lower fluid, respectively. The electrodes are located at $z=\pm L$, where a DC voltage of $2V$ is applied. The position of the flat membrane is $z=0$. Both electrolytes are assumed to be $1:1$. The solutions have the same permittivity $\varepsilon_{f}$, but different concentration $n_{i}^{\infty}$. This implies different conductivities of the electrolyte solutions, $\sigma_{i}=2e^{2}\omega n_{i}^{\infty}$ [14], where $\omega$ is the mobility of ions and $e$ is the elementary charge.
Throughout this study, we assume osmotic equilibration across the membrane, such that no net transmembrane water flux arises from osmotic pressure differences. In experimental systems, this condition can be achieved by adding non-permeating solutes—such as sugars like sucrose or glucose—to ensure equal osmotic pressure on both sides of the membrane [26].
Figure 1: A planar lipid bilayer membrane, with shape denoted by $z=h(x)$, separating electrolyte solutions with different concentrations. The electrodes, with potential $\phi=\pm V$, are held at $z=\pm L$.
@@ II.1 Governing equations
where $\rho_{i}=e(n_{i}^{+}-n_{i}^{-})$ is the charge density, which we assume satisfies the linearized PNP equations
@@
The perturbations in the concentration $n_{i}^{\pm}=n_{i}^{\infty}+\delta n_{i}^{\pm}$ [20] are used to linearize the PNP, assuming small applied voltages compared to thermal voltage $k_{B}T/e$ [27]. $\bm{j}_{i}$ is the electric current density, $k_{B}T$ is the thermal energy, $D$ is the ion diffusion coefficient (all ions are assumed to have the same mobility), and $\bm{v}_{i}$ is the fluid velocity. Here we introduce the Debye length
$\lambda_{D_{i}}=\left(\frac{2e^{2}n^{\infty}_{i}}{\varepsilon_{f}k_{B}T}\right)^{-1/2},$ (3)
which is the characteristic screening length, and we define $\lambda_{D}=(\lambda_{D_{1}}\lambda_{D_{2}})^{1/2}$ as the composite Debye length in our asymmetric system, following the notation of Ref. [28].
@@
where $\mu_{i}$ is the fluid viscosity and $\rho_{i}\nabla\phi_{i}$ is the Coulomb body force due to the electric field acting on the charge distributed in the bulk.
@@ II.2 Boundary conditions
At the top and bottom electrodes $z=\pm L$, we impose a voltage $\phi_{i}(z=\pm L)=\pm V$. We neglect the polarization at the electrode surfaces [20] since we focus on the membrane dynamics. The distance between the electrodes is assumed to be much larger than the Debye length $L\gg\lambda_{D}$, such that the bulk electrolyte is electroneutral and thus charge density at the electrodes vanishes $\rho_{i}(z=\pm L)=0$.
The membrane’s position is represented by $z=h(x)$, with a unit normal vector $\bm{n}=(-h_{x},1)/(1+h_{x}^{2})^{1/2}$ pointing upward, and a unit tangent vector $\bm{\tau}=(1,h_{x})/(1+h_{x}^{2})^{1/2}$. On the membrane, a Robin-type boundary condition [19, 20]
@@
relates the transmembrane potential $V_{m}=\phi_{1}|_{z=h}-\phi_{2}|_{z=h}$ and the electric displacement field. $C_{m}=\varepsilon_{m}/d_{0}$ is the capacitance of the membrane with the thickness $d_{0}$ and permittivity $\varepsilon_{m}$. This boundary condition was derived by mapping the continuity of the potential $\phi_{i}|_{z=h}=\phi_{m}|_{z=\pm d_{0}/2}$ and the continuity of the displacement field $\varepsilon_{f}\bm{n}\cdot\nabla\phi_{i}|_{z=h}=\varepsilon_{m}\bm{n}\cdot\nabla\phi_{m}|_{z=\pm d_{0}/2}$ to a zero-thickness membrane [19], where $\phi_{m}$ is the electric field inside the membrane and $i=1(2)$ corresponds to $(-)d_{0}/2$. The continuity of the displacement field holds under the assumption of no adsorbed charges at the membrane interface. It is important to note that while the charge in the Debye layer at the membrane may be nonzero, the surface charge is zero. In fact, the bulk charge integrated over the Debye layer in the electrokinetic model [28], expressed as $Q=\int_{\lambda_{D}}\rho dz$, corresponds to the surface charge in the leaky dielectric model. For a flat membrane, we calculate the macroscopic charge from Eq. (1):
$Q_{1}=\int_{z=h}^{z=+\infty}\rho_{1}dz\\ =\left.\varepsilon_{f}\frac{\partial\phi_{1}}{\partial z}\right|_{z=h}-\left.\varepsilon_{f}\frac{\partial\phi_{1}}{\partial z}\right|_{z=\infty}\approx\left.\varepsilon_{f}\frac{\partial\phi_{1}}{\partial z}\right|_{z=h},$ (6)
since $z=+\infty$ is taken as the position far from the membrane, and electric field vanishes in the bulk $\frac{\partial\phi_{1}}{\partial z}|_{z=+\infty}=0$. The macroscopic charge on the other side of the membrane, denoted as $Q_{2}$, is calculated similarly. Upon analysis, the Robin condition in Eq. (5) further reveals $Q_{1}=-Q_{2}$, signifying that the total charge on both sides of the membrane always remains balanced, regardless of conductivity asymmetry.
Another boundary condition is the no-flux condition [29]:
@@
This corresponds to an insulating membrane, i.e., charge is not permitted to cross the membrane.
@@
where we denote $f{]\!]}_{a}=f_{1}(z=a)-f_{2}(z=a)$ as the jump of $f$ at position $z=a$. $\bm{T}$ is the total stress, consisting of the hydrodynamic stress $\bm{T}^{hd}$} and Maxwell’s stress $\bm{T}^{el}$}
@@
where $H$ is the curvature, $K_{G}$ is the Gaussian curvature, and $\Gamma$ is the membrane tension. Note that for an incompressible membrane, the tension is determined by the membrane’s mechanical response [11]. The tangential electric stresses are balanced by gradients in the membrane tension, such that non-uniform tension arises to enforce local area incompressibility. In the linearized formulation, however, the tension remains uniform at leading order, and variations appear only at higher order [11]. $C_{0}$ is the bilayer spontaneous curvature arising from the asymmetry of the monolayers. In this stability study, we initially neglect $C_{0}$ by assuming $C_{0}d_{0}\ll 1$ [31] for a membrane of zero thickness. We come back to this effect in Sec. IV.3.
@@
$\frac{\partial h}{\partial t}=v-u\frac{\partial h}{\partial x}.$ (11)
@@ II.3 Dimensionless parameters and rescaling
All variables are scaled using their characteristic values:
@@
$(x,z)$ $\mapsto$ $\lambda_{D}(x,z),\quad h\mapsto\lambda_{D}h,\quad d_{0}\mapsto\lambda_{D}d,$ (12)
$t$ $\mapsto$ $\underbrace{\frac{\mu\lambda_{D}^{2}}{\varepsilon_{f}\phi_{c}^{2}}}_{t_{el}}t,\quad p\mapsto\frac{\mu}{t_{el}}p,$
where $\phi_{c}$ is the thermal voltage, and $t_{el}$ is the flow time scale [10]. Using this time scale, the scaled PNP model takes the form:
@@
where $R=\frac{\sigma_{1}}{\sigma_{2}}=\frac{n_{1}^{\infty}}{n_{2}^{\infty}}$ denotes the conductivity ratio. With $R\mapsto R^{-1}$, Eqs. (13) applies for phase 2. The parameter $\alpha$ represents the ratio between $t_{el}$ and the characteristic charging time of the double-layer capacitor $t_{c}=\frac{\lambda_{D}L}{D}$ [27, 28]:
@@
In this study, we consider the situation where $\alpha\ll 1$, and set $\alpha=0$. In this limit, the charge relaxation in Debye layer happens instantaneously, and the Debye layer remains at equilibrium while the membrane is deforming. The dimensionless Robin condition is given by
@@
is the dimensionless capacitance. The no-flux condition at the membrane retains the form in Eq. (7) with the dimensionless current
@@
$p+2\partial_{n}v_{n}\\ +\frac{1}{2}[(\partial_{n}\phi)^{2}-(\partial_{\tau}\phi)^{2}]-\frac{1}{2}\beta d[(\partial_{n}\phi_{m})^{2}-(\partial_{\tau}\phi_{m})^{2}]{]\!]}_{z=h}\\ =-\frac{1}{Ca}(4H^{3}-4HK_{G}+2\nabla_{s}^{2}H-2\bar{\Gamma}H),$ (19)
where $\partial_{n}v_{n}=\bm{n}\cdot[\bm{\mathrm{\nabla}}\bm{v}+(\bm{\mathrm{\nabla}}\bm{v})^{\top}]\cdot\bm{n}$ is the projection of the viscous stresses, and $\partial_{\tau}\phi=\bm{\mathrm{\nabla}}\phi\cdot\bm{\tau}$. The capillary number $Ca$ compares the electric and bending stresses:
@@
In Sec. III, we will show how the electric field impacts the force balance on the membrane via the Maxwell stress and the hydrodynamic pressure $p$. The traction from the Maxwell stress at the membrane consists of two parts, with $\frac{1}{2}(\partial_{z}\phi)^{2}{]\!]}$ coming from the electric field in the external medium, and $\frac{1}{2}\beta d(\partial_{z}\phi_{m})^{2}{]\!]}$ coming from the electric field $\bm{E}_{m}=-\nabla\phi_{m}$ in the membrane [32, 33]. We emphasize that although the Robin condition in Eq. (15) effectively maps the electric field and charge density on both sides of a finite-thickness membrane to a zero-thickness membrane, the $\phi_{m}$ term in Eq. (19) remains necessary to calculate the effective stress in the zero-thickness limit [17, 20, 18]. We will address this in Sec. III.